OpenStudy (anonymous):

medal&fannnnn A scale drawing of a bathroom is shown below: A rectangle is shown. The length of the rectangle is labeled 2 inches. The width of the rectangle is labeled 3.5 inches. The scale is 1 to 20. What is the area of the actual bathroom in square feet? Round your answer to the nearest whole number. 7 ft to the power of 2 19 ft to the power of 2 12 ft to the power of 2 140 ft to the power of 2

undeadknight26 (undeadknight26):

LOL This was funny for some reason.

OpenStudy (igreen):

?

OpenStudy (igreen):

Well since the scale is 1:20, you have to scale it by 20:1..

OpenStudy (igreen):

20 * 2 = ? 20 * 3.5 = ? @KraziiKai Can you multiply those?

OpenStudy (anonymous):

40&70? @iGreen

OpenStudy (igreen):

Yep, you got it! So the length's gonna be 40, and the width's gonna be 70.

OpenStudy (igreen):

Now we have to find the area.

OpenStudy (igreen):

Area of a rectangle = \(l \cdot w\)

OpenStudy (anonymous):

hmm I'm still confused.. llol @iGreen